# Interval Decomposition of Infinite Persistence Modules over a Principal Ideal Domain

### v1

@@ 4.3. Incremental Extension
If $\textrm{PD}_{n}^{F}(\mathcal{K})$ is identical for every coefficient field $F$, then the same holds for $\mathcal{K}|_{S}$, for every finite subset $S\subseteq I$. It follows by [12, Theorem 1.9] that (4) implies (3). The argument from [10, Theorem 4] carries over without modification to show that (3) implies (2). Our main result, Theorem 1.1, shows that (2) implies (1).
Because $H_{n-1}(\mathcal{K}_{a};\mathbb{Z})$ and $H_{n}(\mathcal{K}_{a};\mathbb{Z})$ are free for all $a$, the maps $(\dagger)$ in (5.1) are isomorphisms, by the universal coefficient theorem. The extension maps ()(*) induce homomorphisms $\varphi:\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})\to\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})\otimes F$ of persistence modules, which preserves direct sums, in the sense that $\varphi(\bigoplus_{i}A_{i})=\bigoplus_{i}\varphi(A_{i})$. Therefore, the same holds for the map $\psi:\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})\to\mathcal{H}_{n}(\mathcal{K};F)$ obtained by composing $\varphi$ with the isomorphism $\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})\otimes F\xrightarrow{\cong}\mathcal{H}_{n}(\mathcal{K};F)$. Consequently, if $\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})$ splits as a direct sum of interval submodules, then $\psi$ yields a corresponding decomposition of $\mathcal{H}_{n}(\mathcal{K};F)$, with the same multiset of intervals. This establishes that (1) implies (4). The same homomorphism carries consistent bases to consistent bases. ∎
@@ Appendix A A useful short exact sequence
To check exactness at $A_{i_{t}}^{pq}$, because $\pi$ is surjective, it suffices show that $f(a\leq i_{t})$ restricts to a surjection $\mathrm{IK}_{a}[i_{p},i_{q}]\to\mathrm{IK}_{i_{t}}[i_{p},i_{q}]$. By [10, Theorem 15], we have $f(a\leq i_{t})(\mathrm{IK}_{a}[i_{p},i_{q}])=\mathrm{IK}_{i_{t}}[i_{p},i_{q}]\cap\mathrm{Im}[a,i_{t}]$. However, because $i_{p}\leq a$, we have $\mathrm{Im}[i_{p},i_{t}]\subseteq\mathrm{Im}[a,i_{t}]$, which implies that
@@
Note that by [10, Theorem 16], we have

### v2

@@ 4.3. Incremental Extension
If $\textrm{PD}_{n}^{F}(\mathcal{K})$ is identical for every coefficient field $F$, then the same holds for $\mathcal{K}|_{S}$, for every finite subset $S\subseteq I$. It follows by [15, Theorem 1.9] that (4) implies (3). The argument from [13, Theorem 4] carries over without modification to show that (3) implies (2). Our main result, Theorem 1.1, shows that (2) implies (1).
Because $H_{n-1}(\mathcal{K}_{a};\mathbb{Z})$ and $H_{n}(\mathcal{K}_{a};\mathbb{Z})$ are free for all $a$, the maps $(\dagger)$ in (5.1) are isomorphisms, by the universal coefficient theorem. The extension maps ()(*) induce homomorphisms $\varphi:H_{n}(\mathcal{K};\mathbb{Z})\to H_{n}(\mathcal{K};\mathbb{Z})\otimes F$ of persistence modules, which preserves direct sums, in the sense that $\varphi(\bigoplus_{i}A_{i})=\bigoplus_{i}\varphi(A_{i})$. Therefore, the same holds for the map $\psi:H_{n}(\mathcal{K};\mathbb{Z})\to H_{n}(\mathcal{K};F)$ obtained by composing $\varphi$ with the isomorphism $H_{n}(\mathcal{K};\mathbb{Z})\otimes F\xrightarrow{\cong}H_{n}(\mathcal{K};F)$. Consequently, if $H_{n}(\mathcal{K};\mathbb{Z})$ splits as a direct sum of interval submodules, then $\psi$ yields a corresponding decomposition of $H_{n}(\mathcal{K};F)$, with the same multiset of intervals. This establishes that (1) implies (4). The same homomorphism carries consistent bases to consistent bases. ∎
@@ Appendix A A useful short exact sequence
To check exactness at $A_{i_{t}}^{pq}$, because $\pi$ is surjective, it suffices show that $f(a\leq i_{t})$ restricts to a surjection $\mathrm{IK}_{a}[i_{p},i_{q}]\to\mathrm{IK}_{i_{t}}[i_{p},i_{q}]$. By [13, Theorem 15], we have $f(a\leq i_{t})(\mathrm{IK}_{a}[i_{p},i_{q}])=\mathrm{IK}_{i_{t}}[i_{p},i_{q}]\cap\mathrm{Im}[a,i_{t}]$. However, because $i_{p}\leq a$, we have $\mathrm{Im}[i_{p},i_{t}]\subseteq\mathrm{Im}[a,i_{t}]$, which implies that
@@
Note that by [13, Theorem 16], we have
@@
Acknowledgements
We thank Marzieh Eidi for helpful conversations. We also thank the anonymous referee for helpful comments and suggestions.